Suppose you want to operate an ideal refrigerator
Answer #1

According to the question we need to find the temp of the hot reservoir.

Cold temp = -20 deg C

Cop = 6.5

COP = (Tc)/(Th - Tc)

6.5 = 253.15 / ( Th - 253.15 ) = 18.946 deg C ==============ANSWER)